# Circuit-tuning: A Mechanistic Approach for Identifying Parameter Redundancy and Fine-tuning Neural Networks

### v2

@@ Abstract
The study of mechanistic interpretability aims to reverse-engineer a model to explain its behaviors. While recent studies have focused on the static mechanism of a certain behavior, the learning dynamics inside a model remain to be explored. In this work, we develop an interpretable fine-tuning method for analyzing the mechanism behind learning. We first introduce the concept of node-level intrinsic dimensionality to describe the learning process of a model in a computational graph. Based on our theory, we propose circuit-tuning, a two-stage algorithm that iteratively builds the minimal subgraph for a specific task and updates the key parameters in a heuristic way. Experimental results confirm the existence of the intrinsic dimensionality at the node level and demonstrate the effectiveness of our method for transparent and interpretable fine-tuning. We visualize and analyze the circuits before, during, and after fine-tuning, providing new insights into the self-organization mechanism of a neural network in the learning process.
@@ 1 Introduction
Transformer-based large language models (LLMs) have demonstrated outstanding performance in a wide range of tasks [61]. However, an LLM is often treated as a "black box” because of its complex inner mechanisms, which brings a lot of issues about AI saftey and reliability, highlighting the need for interpretability [24]. Mechanistic interpretability aims to discover the underlying mechanisms inside a model so as to provide better control and improved design of it [53], showing the potential of reverse-engineering a model. Recent studies in this field include analyzing the circuit responsible for a single behavior [41, 62], extracting features via sparse dictionary learning [5, 57], applying a steering vector [60] to modify the model behaviors, etc.
@@
Despite the success of the above methods, they are limited to post-hoc analyses of a trained model. For example, Wang et al. [62] and Hanna et al. [18] studied the interactions of the attention heads / MLPs and discovered the circuits for indirect object identification and mathematics in GPT-2-small. This kind of static analysis of model behaviors during inference fails to explain the learning dynamics of a model, i.e., how a model acquire an ability, or generalize to various scenarios, which is of great importance in mechanistic interpretability [53]. Recently, Nanda et al. [38], Olsson et al. [42] studied the phase transition during training. Prakash et al. [46], Lee et al. [28], Jain et al. [23], Wang et al. [63] focused on narrow or synthetic tasks for fine-tuning. Bricken et al. [4] studied the change in sleeper agent features via stage-wise model diffing. While these works focus on specific scenerios, there remain two limitations: (1) there lacks a general and unified way or pipeline for interpreting the learning process that is applicable to various scenarios from the view of mechanistic interpretability; (2) current studies mostly focus on post-hoc interpretations of fine-tuning, without daring to provide guidance for a more precise and effective fine-tuning process with interpretability. These limitations make it difficult for interpretability works to be well applied in practice to more general settings, or to provide more inspiration for traceable and steerable training.
To solve the above two concerns, we introduce a general and interpretable fine-tuning method, making it possible for transparent training. Specifically, inspired by prior works on intrinsic dimensionality, we view the learning process inside a model as finding and optimizing the minimum subgraph corresponding to the intrinsic dimensions in a computational graph. Previous estimation of intrinsic dimensionality was conducted directly in the parameter space, while the research on mechanistic interpretability usually adopts a representational perspective. To bridge these two paradigms, we introduce the node-level intrinsic dimensionality for the first time, enabling the study of learning dynamics with intrinsic dimensionality in the representation space.
Based on our theory, we propose circuit-tuning, a method that performs interpretable fine-tuning in a computational graph. Our method is able to: (1) precisely and dynamically localize the key parts in a computational graph during fine-tuning on any tasks and (2) construct and optimize the circuit to fit a new data distribution, thereby actively providing guidance for the fine-tuning process from an interpretable perspective. Experiments confirm the validity of our theory and the effectiveness of our method on various scenarios. We visualize and analyze a transparent, interpretable, and traceable training process, provide new findings about the self-organization mechanism inside a model during fine-tuning, and open new horizons for controllable and steerable training.
In summary, our contributions are mainly as follows:
  • We redefine intrinsic dimensionality and extend the estimation of it from the parameter space to the representation space, thereby introducing that concept into mechanistic interpretability and making it possible to study the learning dynamics in a computational graph.
  • We propose circuit-tuning, a general way of interpretable fine-tuning. Our method integrates heuristic and dynamic circuit discovery into fine-tuning for the study of learning dynamics for the first time, serving as a feasible way of interpretability-driven training.
  • We confirm the effectiveness of our method through experiments and gain new insights into the mechanisms behind learning through causal analysis and visualization, providing new insights and improvements for the study of learning dynamics.
@@ 2 Related work
2.1 Intrinsic dimensionality
Intrinsic dimension was studied in early works to estimate the dimensionality of a dataset [15, 58, 6, 29]. Inspired by the redundancy in parameters, Li et al. [31] defined intrinsic dimensionality as the minimum number of parameters needed to reach satisfactory solutions for an objective function. Aghajanyan et al. [1] focused on fine-tuning and defined it as the minimal description length in terms of a specific task, providing new insights into the generalization during pre-training. Prior works directly looked at the weights in a model and measure the intrinsic dimensionality in the parameter space, while the decomposition of a model and the following research on mechanistic interpretability focused on the activations in the representation space. The divergence of research perspectives has brought difficulties to the study of intrinsic dimensionality in the field of mechanistic interpretability.
2.2 Mechanistic interpretability for learning
@@
Mechanistic interpretability aims to understand the computational mechanisms of a model [53]. Existing works can be divided into static research on the trained model and dynamic research during training according to whether the model is in an inference state or a training state. In the static research, circuit discovery is a widely used technique that aims to look for the subgraph in a computational graph that acts as an algorithm implemented in a model for a certain behavior. Meng et al. [35] proposed activation patching to identify activations relevant to the output, while Nanda [37] proposed attribution patching to accelerate it. Wang et al. [62], Conmy et al. [9] focused on edges and proposed path patching. Others optimized this technique from various aspects [56, 27, 34, 2, 32]. Despite their success, current studies are limited to discover the circuit of an existing model behavior of a trained model in a static style, while our method is able to form a new circuit of a non-existent capability through iterations in a dynamic mode during training.
@@
Compared to static research, there is fewer research on learning dynamics. For phase transition, Nanda et al. [38] delved into the study of grokking, while Olsson et al. [42] analyzed the emergency of induction heads for in-context learning. For fine-tuning mechanisms, Prakash et al. [46] focused on entity linking, Jain et al. [23] focused on compiled models and probabilistic context-free grammars, Wang et al. [63] focused on math, while Lee et al. [28] focused on DPO for toxicity reduction. Kotha et al. [26] explored the catastrophic forgetting. Wu et al. [64] proposed ReFT that learns an intervention on the representations for efficient fine-tuning. Parallel to our work, Ren and Sutherland [50] proposed a framework for learning dynamics by decomposing the change of a model’s prediction. However, current studied are limited to a single scenario with various analytic methods, while there lacks a unified method for the study of learning dynamics on common tasks, and no attempt was made to actively introduce interpretability into fine-tuning for mechanistic study.
Unlike prior works, we successfully transfer the concept of intrinsic dimensionality from the parameter space to the representation space, and dynamically integrate circuit discovery into fine-tuning as a heuristic method for the study of learning dynamics for the first time, providing new insights for the study of learning dynamics from the mechanistic view.
3 The theory of node-level intrinsic dimensionality
3.1 Background
@@
As discussed in Section 2.1, intrinsic dimensionality implies the redundancy in parameters. Measuring the exact intrinsic dimensionality is computational intractable. Therefore heuristic methods are adopted to measure the upperbound of it. Considering a set of parameters $\theta^{D}=[\theta_{0},...,\theta_{m}]$ and an objective function $f(\cdot,\theta)$, Li et al. [31] leveraged a re-parameterization method to optimize only the parameters $\theta^{d}$ in a subspace $\mathbb{R}^{d}(d<D)$ via a linear transformation with a projection matrix $P$:
@@
If a satisfactory solution is reached, then the dimensionality of that subspace is the intrinsic dimensionality. In practice, the heuristic method requires searching over various $d$, optimizing the parameters $\theta^{d}$ and selecting the smallest $d$ that could reach a satisfactory result.
3.2 Redefine intrinsic dimensionality
In this section, we view the model $M$ as a computational graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ where $\mathcal{V}$ and $\mathcal{E}$ represent the nodes and edges in $\mathcal{G}$ respectively. Each node is a vector $n=(h_{1},...,h_{N})^{\top}\in\mathbb{R}^{N}(1\leq N\leq D)$ that could be the activation of a neuron, a group of neurons or a representation in a $D$-dimensional representation space $\mathbb{V}^{D}$, based on granularities of interest. The edges describe the information flows between the nodes, i.e., where the output of an upstream node goes and where the input of a downstream node is from.
Note that it is possible to set a node to a full layer, or a latent which is a sparse feature in dictionary learning, while here we require the size of a node to be no more than the dimensionality of $\mathbb{V}^{D}$ to cover most of the cases in practice and for convenience of discussion. Also, an edge is not necessary to follow the real structure of a model. It could be a virtual connection between non-adjacent nodes.
Since the decomposition of a model in mechanistic interpretability often looks at activations or latents [53], we look for the intrinsic dimensions indirectly via the activations in the representation space. Inspired by prior works on circuit discovery, we believe that the learning process of a model is to dynamically identify the redundancy in a computational graph and optimize the subgraph of a task.
Definition 1: intrinsic dimensionality
Given the computational graph $\mathcal{G}$ of a model and a task $T$, the intrinsic dimensionality is the number of nodes in the minimal subgraph $\mathcal{C}^{*}=\{\mathcal{V}_{T}^{*},\mathcal{E}_{T}^{*}\}\subset\mathcal{G}$ on which the optimization of the objective function for task $T$ could reach a predefined precision level.
The newly defined intrinsic dimensionality, namely the node-level dimensionality, describes the redundancy in parameters indirectly with the "intrinsic nodes" in the minimal subgraph. Under this new definition, the parameters necessary for a task $T$ refer to those where the intrinsic nodes $\mathcal{V}_{T}^{*}$ are derived from. For example, if a parameter matrix $W\in\mathbb{R}^{m\times n}$ maps an input $x\in\mathbb{R}^{n}$ to a representation $H\in\mathbb{R}^{m}$ that is an intrinsic node, then $W$ is what matters to task $T$. A theoretical explanation of this is shown in Section 3.4 and detailed in Appendix E.
3.3 Estimating intrinsic dimensionality
Different from prior works that measure the intrinsic dimensionality in a subspace, we choose to perform the estimation in a subgraph. Similar to Li et al. [31], we also adopt a heuristic method to estimate the intrinsic dimensionality through training. The difference is that instead of optimizing via re-parameterization, we directly search for the intrinsic nodes in the computational graph through iterations. Inspired by the idea of causal intervention for the structural causal model [39], we adopt the concept of indirect effect [44]. Specifically, let’s consider a node $n\in\mathcal{V}$. For the task $T$, a batch of inputs $\mathcal{X}=\{x_{i}\}_{i=1}^{p}\sim\mathcal{D}_{T}$ is fed into the model $M$. For each input $x$, if we replace $n(x)$ that is the activation of the node $n$ with another value $n(x^{\prime})$ from a corrupted input $x^{\prime}$ while keeping other activations unchanged, we can get another output $M(x^{\prime})$. Given a metric $\mathcal{L}_{m}$ for measuring the output, the contribution $c(n)$ of the node $n$ in terms of the input $x$ is its indirect effect on the output:
$c(n)=\Big{|}\mathcal{L}_{m}\big{[}M\big{(}x\,|\,do\big{(}n\leftarrow n(x^{\prime})\big{)}\big{)}\big{]}-\mathcal{L}_{m}\big{[}M(x)\big{]}\Big{|}$ (2)
The do-calculus notation [39] is used to express the intervention behavior in Equation (2). Given a lowerbound $\tau$, If the contribution of $n$ satisfies: $\mathbb{E}_{x_{i}\sim\mathcal{X}}\big{[}c_{i}(n)\big{]}>\tau$, then $n$ is an intrinsic node. During optimization, we dynamically prune the graph to encourage exploration over the computational graph. When a satisfactory precision of optimization is reached, then the intrinsic dimensionality is the minimum number of intrinsic nodes $|\mathcal{V}_{T}^{*}|$ in the final subgraph for the task $T$.
3.4 Further discussions on the node-level intrinsic dimensionality
What does the indirect effect imply?
As discussed in Section 3.3, the indirect effect is adopted to measure the importance of a node in a causal manner. Since intrinsic dimensionality originates from the parameter redundancy in a model [29, 1], while a node is defined over the activations, how can we describe parameter redundancy through the lens of activations, and what exactly are the redundancies in a representation? To gain a deeper understanding of it, we provide a theoretical analysis detailed in Appendix E. In short, we decompose Equation (2) into two parts by applying a first-order Taylor expansion to the right side of it at $n=n(x)$ for approximation:
(2$)\!\approx\!\Big{|}\mathcal{L}_{m}[M(x)]\!+\![n(x^{\prime})\!-\!n(x)]^{\top}\nabla_{n}\mathcal{L}_{m}[M(x)]\!-\!\mathcal{L}_{m}[M(x)]\Big{|}\!=\!\Big{|}[n(x)\!-\!n(x^{\prime})]^{\top}\!\nabla_{n}\mathcal{L}_{m}[M(x)]\Big{|}$ (3)
In practice, the intervention on the node $n$ is performed using the averaged result of all other nodes. Therefore $n(x)-n(x^{\prime})$ can be viewed as a centered activation $n(x)^{c}$. Thus Equation (3) consists of two parts: (i) the first part $n(x)^{c}$ is an indicator for representation redundancy, which describes the redundancy of a node for representing an input; (ii) the second part $\nabla_{n}\mathcal{L}_{m}[M(x)]$ is an indicator for forward redundancy, which describes the redundancy of a node in the forward propagation. Thus the estimation of the intrinsic dimensionality can be converted to the process of redundancy detection in a graph. The definitions of the redundancies as well as the proof are detailed in Appendix E.
The minimum description length (MDL) in a computational graph.
The MDL principle [51] encourages us to minimize the combined cost of encoding the model and the data given the model. While the parameters for the task $T$ may not be decoupled from others and are perhaps distributed over the whole parameter space $\mathbb{R}^{D}$, prior works on intrinsic dimensionality used a reparameterization trick to separate the intrinsic dimensions $\theta^{d}\in\mathbb{R}^{d}$ out from the original parameter space. This kind of measurement aims to look for the upperbound of the MDL in terms of a task.
While we aim to (i) find out the minimal subgraph $\mathcal{C}^{*}$ from a computational graph $\mathcal{G}$ and (ii) interpret the learning process of a model based on the subgraph, there is a need to balance the minimum of the required parameters and the convenience for interpretable graph structure. Since we want to keep the original topological structure of the computational graph for further analyses, we directly delve into the original graph to inspect the redundancy, instead of adding extra structures like linear projections. Therefore, the minimal subgraph does not mean the minimum parameters in a strict sense.
However, the functionalities of the nodes in a computational graph are thought to be decoupled from each other in the mechanistic study. Braun et al. [3] argued that one of the properties of a good decomposition of a model is the minimality that the decomposition should use as few components as possible to represent a model behavior of interest, which also follows the MDL principle. Therefore, our definition holds for the MDL principle in the sense of the minimality in mechanistic interpretability. Different from prior works, we do not require a strict estimation of the intrinsic dimensionality, but trying to take into account both the minimality of the graph and the feasibility for interpretability.
4 Main method
4.1 Circuit-tuning
From nodes to edges Following the discussions in Section 3, we introduce circuit-tuning, an interpretable algorithm that allows transparent and precise fine-tuning. The only difference is that we shift our view from nodes to edges. The reason is that (i) we are interested in the interactions between nodes and (ii) the redundancy in nodes can be identified indirectly by checking the contributions of the edges connected to it. The shift from edges to nodes is actually an extension of the estimation method in Section 3.3, and the underlying idea is the same. Thus we can leverage the path patching technique in circuit discovery to identify the intrinsic dimensions. Discussions are detailed in Appendix D.
Given a model $M$ for fine-tuning, the model is initialized into a computational graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, as defined in Section 3.2. A metric $\mathcal{L}_{m}$ and a loss function $\mathcal{L}$ are used for output quantification and parameter optimization respectively, as described in Section 3.3. Given a dataset $\mathcal{X}\sim\mathcal{D}_{T}$ for the fine-tuning task $T$, circuit-tuning alternately performs the following two procedures:
(1) Graph pruning: For a batch of data $\mathcal{X}_{T}\in\mathcal{X}$, we compute the edge contribution for each edge $e:n_{1}\rightarrow n_{2}\in\mathcal{E}$ according to Equation (D), where $n_{1}$ and $n_{2}$ are the upstream and downstream nodes respectively. Then all the edges are sorted in descending order based on their contributions, and the edges with top $N$ contributions are selected. Then the graph $\mathcal{G}$ is pruned into a circuit $\mathcal{C}=\{\mathcal{V}_{T},\mathcal{E}_{T}\}$ with only the selected edges $\mathcal{E}_{T}$ together with the nodes $\mathcal{V}_{T}$ at both ends of each edge inside.
(2) Circuit fine-tuning: For a batch of data $\mathcal{X}_{T}\in\mathcal{X}$, all the parameters outside $\mathcal{C}$ are frozen, and only the parameters corresponding to the nodes $\mathcal{V}_{T}$ (see Section 3.2) are updated through a forward pass and a backward pass as usual. After $K$ steps of optimization, all the frozen parameters are freed and the graph $\mathcal{G}$ is reset to its original state.
In practice, it is efficient to accelerate graph pruning with only one forward and backward pass using edge attribution patching [56]. The implementation details are shown in Algorithm 1 and Section 5.
Algorithm 1 Circuit-tuning
@@ 4.1 Circuit-tuning
  Input: dataset $\mathcal{X}$, model $M$, loss function $\mathcal{L}$, the metric for the quantification of the output $\mathcal{L}_{m}$, the number of edges to save $N$, the optimization steps after circuit discovery $K$.
@@
     if $i\bmod K==0$ then
@@
        for edge $e:n_{1}\rightarrow n_{2}\in\mathcal{E}$ do
           Intervene in edge $e$ and calculate the contribution of it according to Equation (7) and (D)
        end for
[…]
@@
     Update the parameters corresponding to the nodes in $\mathcal{C}$
@@
4.2 Characteristics of Our Method
Heuristic and dynamic circuit discovery Unlike prior works that focus on the circuits in a trained model, circuit-tuning acts as a heuristic approach for a model to develop an unseen ability in a dynamic style. The validity of this idea is built on the finding in [1] that pre-training has learned enough knowledge for downstream tasks, allowing fine-tuning with minor modifications.
Interpretable model diffing Instead of being a blackbox, circuit-tuning performs fine-tuning on a visible circuit, making the full process open to researchers. The transparency of our method provides the convenience to track the learning dynamics and the availability to intervene in the training process. Thus we are able to compare among different training stages and gain a deeper understanding and better control of it. Visualization of the training process is detailed in Section 5 and Appendix F.4.
Another feature of our method is that it is able to precisely locate the key components and update only a few parameters as discussed in Section 3.4, instead of modifying all of the parameters as LoRA-based methods do. Besides, while ReFT [64] a similar method also leverages representations for training, it is aimed at parameter-efficient fine-tuning and learns a transformation as an intervention on a hidden state, while our method is more focused on the mechanisms of learning during training.
5 Experiments
In this section, we test our theory and method across a variety of models and tasks. To comprehensively investigate the learning dynamics during fine-tuning as well as the practicality of our method, we arrange the experiments into two parts. The first part serves as the verification and analysis of our method, while the second part aims to show that our method is applicable to various scenarios.

### v3

# Fine-Tuning is Subgraph Search: A New Lens on Learning Dynamics

@@ Abstract
The study of mechanistic interpretability aims to reverse-engineer a model to explain its behaviors. While recent studies have focused on the static mechanism of a certain behavior, the learning dynamics inside a model remain to be explored. In this work, we develop a fine-tuning method for analyzing the mechanism behind learning. Inspired by the concept of intrinsic dimension, we view a model as a computational graph with redundancy for a specific task, and treat the fine-tuning process as a search for and optimization of a subgraph within this graph. Based on this hypothesis, we propose circuit-tuning, an algorithm that iteratively builds the subgraph for a specific task and updates the relevant parameters in a heuristic way. We first validate our hypothesis through a carefully designed experiment and provide a detailed analysis of the learning dynamics during fine-tuning. Subsequently, we conduct experiments on more complex tasks, demonstrating that circuit-tuning could strike a balance between the performance on the target task and the general capabilities. Our work offers a new analytical method for the dynamics of fine-tuning, provides new findings on the mechanisms behind the training process, and inspires the design of superior algorithms for the training of neural networks. Code is available at https://github.com/Siriuslala/circuit-tuning.
@@ 1 Introduction
Transformer-based large language models (LLMs) have demonstrated outstanding performance in a wide range of tasks (Vaswani, 2017). However, an LLM is often treated as a ”black box” because of its complex inner mechanisms, which brings a lot of issues about AI saftey and reliability, highlighting the need for interpretability (Ji et al., 2023). Mechanistic interpretability aims to discover the underlying mechanisms inside a model so as to provide better control and improved design of it Sharkey et al. (2025), showing the potential of reverse-engineering a model. Recent studies in this field include analyzing the circuit responsible for a single behavior Olah et al. (2020); Wang et al. (2022), extracting features via sparse dictionary learning (Bricken et al., 2023; Templeton et al., 2024), applying a steering vector (Turner et al., 2023) to modify the model behaviors, etc.
@@
Despite the success of the above methods, they are limited to post-hoc analyses of a trained model. For example, Wang et al. (2022) and Hanna et al. (2023) studied the interactions of the attention heads / MLPs and discovered the circuits for indirect object identification and mathematics in GPT-2-small. This kind of static analysis of model behaviors during inference fails to explain the learning dynamics of a model, i.e., how a model acquire an ability, or generalize to various scenarios, which is of great importance in mechanistic interpretability (Sharkey et al., 2025). Recently, Nanda et al. (2023); Olsson et al. (2022) studied the phase transition during training. Prakash et al. (2024); Lee et al. (2024); Jain et al. (2023); Wang et al. (2025) focused on narrow or synthetic tasks for fine-tuning. Bricken et al. (2024) studied the change in sleeper agent features via stage-wise model diffing. While these works focus on specific scenarios with various analytic methods, there still remain two limitations: (1) there lacks a general and unified way or pipeline for interpreting the learning process that is applicable to various scenarios from the view of mechanistic interpretability; (2) current studies mostly focus on post-hoc interpretations of fine-tuning, without daring to provide guidance for a more precise and effective fine-tuning process with interpretability. These limitations make it difficult for interpretability works to be well applied in practice to more general settings, or to provide more inspiration for traceable and steerable training.
Figure 1: The overall pipeline of our work. Circuit-tuning iteratively builds the subgraph for a specific task and updates the relevant parameters. The intermediate subgraphs are saved during fine-tuning, and various interpretability techniques could be utilized for the study of learning dynamics.
To solve the above two concerns, we introduce a intuitive and heuristic fine-tuning method, making it possible for transparent training. Specifically, we view the model as a computational graph, where the nodes are terms in its forward pass (neurons, attention heads, etc.) and the edges are the interactions between the nodes. Inspired by prior work on intrinsic dimension—which suggests that only a small subset of dimensions in a neural network is useful when fine-tuning on a specific task—we argue that for the fine-tuning process on a given task, the computational graph of the model is similarly redundant, which could lead to unnecessary updates to the graph. We believe that the learning and generalization process of the model on a specific task is to find and optimize the subgraph in a computational graph. Following this idea, we propose circuit-tuning, a method that performs fine-tuning in a computational graph. Our method is able to: (1) precisely and dynamically localize the key parts in a computational graph during fine-tuning on any tasks and (2) construct and optimize the circuit to fit a new data distribution, thereby actively providing guidance for the fine-tuning process from an interpretable perspective.
To test the effectiveness of our method, we firstly designed an experiment called “subject-verb disagreement” and conducted a study on GPT-2. This experiment is sufficiently small and interesting to allow us to fully validate our hypothesis and method. Through it, we discovered several phenomena during fine-tuning: the functional reversal of specific graph nodes, the preservation of nodes for general capabilities, and the strengthening or weakening of connections between related nodes, and so on. These findings offer a deeper understanding of how fine-tuning works by looking at it through the lens of a computational graph. Subsequently, we performed experiments on more complex tasks and found that circuit-tuning leads to more precise fine-tuning, which improves the performance of downstream tasks while better maintaining general capabilities. This proves that our method can scale to larger models and more complex tasks. It further confirms that circuit-tuning can accurately locate the key model components responsible for a specific task in a dynamic way during training.
In summary, our work provides a new method for studying learning dynamics from the perspective of mechanistic interpretability, and provides a deeper understanding of the fine-tuning process in a computational graph. Our method also provides inspiration for designing better training algorithms that could strike a balance between the performance on the target task and the general capabilities.
@@ 2 Related work
Circuit analysis Mechanistic interpretability aims to understand the computational mechanisms of a model (Sharkey et al., 2025). Existing works can be divided into static research on the trained model and dynamic research during training according to whether the model is in an inference state or a training state. In the static research, circuit analysis is a widely used technique that aims to find and study the subgraph in a computational graph that acts as an algorithm implemented in a model for a certain behavior. Recent studies generally use causal intervention for circuit discovery. Meng et al. (2023) proposed activation patching to identify activations relevant to the output, while Nanda (2023) proposed attribution patching to accelerate it. Wang et al. (2022); Conmy et al. (2023) focused on edges and proposed path patching. Others optimized this technique from various aspects (Syed et al., 2023; Kramár et al., 2024; Marks et al., 2024; Ameisen et al., 2025; Lindsey et al., 2025). Despite their success, current studies are limited to discover the circuit of an existing model behavior of a trained model in a static style, while our method is able to form a new circuit of a non-existent capability through iterations in a dynamic mode during training.
@@
Mechanistic interpretability for learning Compared to the static research mentioned above, there is fewer research on learning dynamics. For phase transition, Nanda et al. (2023) delved into the study of grokking, while Olsson et al. (2022) analyzed the emergency of induction heads for in-context learning. For fine-tuning mechanisms, Prakash et al. (2024) focused on entity linking, Jain et al. (2023) focused on compiled models and probabilistic context-free grammars, Wang et al. (2025) focused on math, while Lee et al. (2024) focused on DPO for toxicity reduction. Kotha et al. (2023) explored the catastrophic forgetting. Wu et al. (2024) proposed ReFT that learns an intervention on the representations for efficient fine-tuning. Parallel to our work, Ren & Sutherland (2024) proposed a framework for learning dynamics by decomposing the change of a model’s prediction. However, current studied are limited to a single scenario, while there lacks a unified method for the study of learning dynamics on common tasks, and no attempt was made to actively introduce interpretability into fine-tuning for mechanistic study. Unlike prior works, we dynamically integrate circuit discovery into fine-tuning as a heuristic method for the study of learning dynamics for the first time, providing new insights for the study of learning dynamics from a mechanistic view.
3 Main method
3.1 Describe the learning process in a mechanistic view
From the view of mechanistic interpretability, we view a model $M$ as a computational graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ which is a directed acyclic graph (DAG), where $\mathcal{V}$ and $\mathcal{E}$ represent the nodes and edges in $\mathcal{G}$, respectively. Each node is a vector $n=(h_{1},...,h_{N})^{\top}\in\mathbb{R}^{N}(1\leq N\leq D)$ that could be the activation of a neuron, a group of neurons or a representation in a $D$-dimensional representation space $\mathbb{V}^{D}$, based on granularities of interest. The edges describe the information flow between the nodes, i.e., where the output of an upstream node goes and where the input of a downstream node is from  [^1] It is possible to set a node to a full layer, or a latent which is a sparse feature in dictionary learning, while here we require the size of a node to be no more than the dimensionality of $\mathbb{V}^{D}$ to cover most of the cases in practice and for convenience of discussion.. An edge is not necessary to follow the real structure of a model, i.e., it could be a virtual connection between non-adjacent nodes.
@@ 3.1 Describe the learning process in a mechanistic view
Given a specific task for training, the parameters in a model is often redundant. Li et al. (2018) defined intrinsic dimensionality as the minimum number of parameters needed to reach satisfactory solutions for an objective function. Considering a set of parameters $\theta^{D}=[\theta_{0},...,\theta_{m}]$ and an objective function $f(\cdot,\theta)$, they adopted a heuristic method to measure the upperbound of the intrinsic dimensionality. They leveraged a re-parameterization method to optimize only the parameters $\theta^{d}$ in a subspace $\mathbb{R}^{d}(d<D)$ via a linear transformation with a pre-defined projection matrix $P$:
@@
If a satisfactory solution is reached, then the dimensionality of that subspace is the intrinsic dimensionality. In practice, the heuristic method requires searching over various $d$, optimizing the parameters $\theta^{d}$ and selecting the smallest $d$ that could reach a satisfactory result. Drawing inspiration from the concept of intrinsic dimension, we make an analogy: given a specific task $T$, redundancy exists in the graph $\mathcal{G}$, and the initial subgraph (circuit) related to this task is $\mathcal{C}=\{\mathcal{V}_{T},\mathcal{E}_{T}\}\subset\mathcal{G}$. Then we propose a hypothesis that during training, the learning process could be described as follows:
The learning process of a model on a task $T$ during training is to dynamically search a subgraph $\mathcal{C}_{i}=\{\mathcal{V}_{T,i},\mathcal{E}_{T,i}\}\subset\mathcal{G}$ for the task at each training step $i$ and adjust the corresponding parameters.
Here the parameters for a subgraph $\mathcal{C}=\{\mathcal{V}_{T},\mathcal{E}_{T}\}$ refer to those where the nodes $\mathcal{V}_{T}$ are derived from. For example, if a parameter matrix $W\in\mathbb{R}^{m\times n}$ maps an input $x\in\mathbb{R}^{n}$ to an activation $H\in\mathbb{R}^{m}$ that is a node in $\mathcal{C}$, then $W$ is what matters to task $T$ (see more examples in Table 2 and 8). In practice, the minimal subgraph $\mathcal{C}^{*}=\{\mathcal{V}_{T}^{*},\mathcal{E}_{T}^{*}\}\subset\mathcal{G}$ is the subgraph when all the redundant components in graph $\mathcal{G}$ are just pruned, which is called an “intrinsic graph”. By removing redundancy from the computational graph, we can concentrate the optimization process on the parameters responsible for the target task. This intuitive approach aims to minimally affect parameters corresponding to functions that are either task-irrelevant or remain invariant to fine-tuning, a process that will be implemented in Section 3.2. The hypothesis for learning will be verified in Section 4.
Algorithm 1 Circuit-tuning (based on edge attribution patching)
@@
Input: dataset $\mathcal{X}$, model $M$, the number of edges to save NN, the number of optimization steps after graph pruning $K$.
@@
  if $i\bmod K==0$ then $\triangleright$ Graph pruning
@@
   Get the edge contribution for each edge via edge attribution patching.
   $\mathcal{E}_{T}\leftarrow\{e\,|\,e\!\in\!\mathcal{E}\text{ with top-$N$ edge contributions}\}$
@@
  Update the parameters corresponding to the subgraph $\mathcal{C}$ $\triangleright$ Circuit fine-tuning
@@
3.2 Circuit-tuning
Following the discussions in Section 3.1, we introduce circuit-tuning, an algorithm that allows transparent and precise fine-tuning. Given a model $M$ and a dataset $\mathcal{X}\sim\mathcal{D}_{T}$ for the fine-tuning task $T$, the model is firstly initialized into a computational graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$. Next, circuit-tuning alternately performs the following two procedures:
(1) Graph pruning For a batch of data $\mathcal{X}_{T}\in\mathcal{X}$, we perform circuit discovery on graph $\mathcal{G}$ for the task $T$. The graph $\mathcal{G}$ is pruned into a circuit $\mathcal{C}=\{\mathcal{V}_{T},\mathcal{E}_{T}\}$ with only the selected edges $\mathcal{E}_{T}$ together with the nodes $\mathcal{V}_{T}$ at both ends of each edge inside.
(2) Circuit fine-tuning Right after (1), all the parameters outside $\mathcal{C}$ are frozen, and only the parameters corresponding to the nodes $\mathcal{V}_{T}$ (see Section 3.1) are updated on a batch of data $\mathcal{X}_{T}$. After $K$ steps of optimization, the frozen parameters are freed and the graph $\mathcal{G}$ is reset to its original state.
The full process is shown in Algorithm 1. For graph pruning, it is generally based on the idea of causal intervention. For each node or edge in the graph, we corrupt it and measure the change in the model’s prediction. The prediction is quantified via a metric $\mathcal{L}_{m}$. The change is regarded as the contribution of that node or edge to the model prediction. All the nodes/edges are sorted in descending order based on their contributions, and those with top $N$ contributions are selected in the subgraph. In practice, we use edge attribution patching (Syed et al., 2023) to accelerate this process. It requires only one forward and backward pass. The details are presented in Appendix C.
During training, only a small part of the parameters are updated, and it is convenient to save the intermediate subgraphs to trace the state of the internals in a model. When the fine-tuning is done, various tools for circuit analysis could be utilized to study the learning dynamics based on the subgraphs, as shown in Figure 1. This makes it possible for model diffing, i.e., we are able to compare among different training stages and gain a deeper understanding of this process, instead of fine-tuning in a “black box”. Besides, unlike prior works that focus on the circuits in a trained model, circuit-tuning acts as a heuristic approach for a model to develop an unseen ability in a dynamic style. The validity of this idea is built on the finding from Aghajanyan et al. (2020) that pre-training has learned enough knowledge for downstream tasks, allowing fine-tuning with minor modifications on a relatively fixed set of parameters. Thus, the subgraphs during fine-tuning are supposed to be highly overlapping, which will be verified through experiments in Figure 11.
4 Analyzing learning dynamics via circuit-tuning